DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, June 24, 2004 Padelford C-36 1:30 pm Kohn-Rossi cohomology for hypersurface isolated singularities

Takao Akahori (Himeji Institute of Technology, Japan)

Let V be a hypersurface isolated singularity in ℂ^{n+1}. This means that

V = {z ∈ ℂ^{n+1} : f(z) = 0}

for some holomorphic function f that has the origin as an isolated critical point. Let M be the following real hypersurface in V:

M = S_{1}^{2n−1}(0) ∩ V = {z ∈ ℂ^{n+1} : f(z) = 0, |z|^{2} = 1}.

Then M is a CR manifold of hypersurface type. Here we give a simple calculation for the Kohn-Rossi cohomology groups of M in this special case.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.